In 12 years, a man will be three times as old as his son; the difference of their present ages is 30, find their present ages
Answer by lwsshak3(11628)

In 12 years, a man will be three times as old as his son; the difference of their present ages is 30, find their present ages
***
at present:
let x=son's age
30+x=father's age
...
12 years hence:
x+12=son's age
30+x+12=42+x=father's age
...
42+x=3(x+12)
42+x=3x+36
2x=6
x=3
30+x=33
...
son's present age=3
father's present age=33